OpenStudy (anonymous):

A rectangular playground is to be fenced off and divided in two by another fence parallel to one side of the playground. 488 feet of fencing is used. Find the dimensions of the playground that maximize the total enclosed area. The shorter side of the playground is ? ft. The longer side of the playground is ? ft. What is the maximum area?

OpenStudy (qwertty123):



OpenStudy (qwertty123):

2L+3w = 488

OpenStudy (qwertty123):

what's Area?

OpenStudy (anonymous):

A=L*W so I have A=W(244-1.5w) ---> A=-1.5w^2+244W

OpenStudy (anonymous):

worthless...thanks for the assist